what volume of water should be added to 500 ml of 2N(f=0.98) NA2CO3 to make it exactly decinormal?
1
Expert's answer
2020-09-08T05:58:45-0400

Ans

Let the initial concentration of Na2CO3 solution be N₁ and volume of solution V₁ and final concentration be N₂ and volume of solution V₂

Given,

N₁ = 2 (factor 0.98)

N₂ = 1/12 (deci-normal)

V₁ = 500 ml

V₂ = ?

We know that,

N₁ x V₁ = N2 x V2

0.98 x 2 x 500 = (1/12) x V2

V2 = 11760 ml

Volume of water to be added is (11760-500) = 11260 ml (Ans.)